A mixture of 1.6 mol of PCl5, 0.92 mol of Cl2 and 0.78 mol of PCl3 is introduced into a 20.0 L reaction vessel at 500K. At this temperature, the equilibrium constant Kc for the reaction: PCl5 (g) = PCl3 (g) + Cl2 (g) is 1.8 × 10-3. Is the reaction mixture at equilibrium? If not what is the direction of the net reaction?


1
Expert's answer
2021-01-18T01:53:21-0500

Concentration of Pcl5=mol/volume

                                       =1.6/20=0.08M

Similarly,.      Cl2=0.92/20

                              =0.046M

               And Pcl3=0.78/20=0.039M

is a reaction quotient,so

Qc=0.039*0.046/0.08

    =0.09435

And from the question,Kc=0.0018

So,Qc>Kc ,the reaction is not equilibrium.

The reactions shifts to the left, towards the reactants. Therefore, Qc> Kc and the reactions shifts towards the reactants.